A COAXIALLY ORIENTED BEAM-RESONATOR ARRANGEMENT FOURIER TRANSFORM MI-CROWAVE (COBRA-FTMW) SPECTROMETER: GOING CRYOGENIC

\underline{JENS-UWE GRABOW}, SAMUEL PALMER, PATRICK THADDEUS, Radio and Geoastronomie Division, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 and Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138.

After the initial experiments of microwave spectroscopy in the time domain ^{a} it took more than two decades before the technique was re-born. First as a steady gas-waveguide^{b}, later as a molecular beam-resonator^{c} application, the experiment became an established spectroscopic method. During the past years a number of improvements were introduced to the technique. Namely the coaxially oriented beam-resonator arrangement (COBRA) ^{d} ^{e} has dramatically improved the resolution and the sensitivity of the Fourier transform microwave (FTMW) spectrometer. Our current efforts are aiming at an improvement of the COBRA-FTMW sensitivity by means of reduction of the thermal noise background, i.e. reducing the 300K thermal noise power of P_{N} = kTRB at room temperature to the equivalent of 77K - the temperature of liquid nitrogen. We will present a detailed theoretical background which is needed to approach the expected gain in S/N for a spectrometer operated at temperatures significantly below the thermal environment.

aR. H. Dicke and R. W. Romer, Rev. Sci. Instrum. 26, 915(1955).

bJ. Ekkers and W. H. Flygare, Rev. Sci. Instrum. 47, 448(1996).

cT. J. Balle and W. H. Flygare, Rev. Sci. Instrum. 52, 33(1981).

dJ. U. Grabow and W. Stahl, Z. Naturforsch. 45a, 1043(1990).

eJ. U. Grabow, W. Stahl, and H. Dreizler, Rev. Sci. Instrum. 67, 4072(1996).